OpenStudy (anonymous):

A(n) =-5 x 2^n-1 , find the fourth and eighth terms of the sequence.

Parth (parthkohli):

1)Substitute n with 4. 2) Substitute n with 8. You get the answers ;)

Parth (parthkohli):

n means the nth term and we need 4th and 8th terms

OpenStudy (anonymous):

A(4)=-5*2^4-1=-81 A(8)=-5*2^8-1=-1281